soybean meal is 14% protein, corn meal is 7% protein, How many pounds of each ahould be mixed together in order to get 280lb mixture with 10% protein.

I have went through a million of these examples in my math lab with no luck can you please help me??

Answer by rfer(16322)

0.14x+0.07(280-x)=0.1(280)
0.14x+19.6-0.07x=28
0.07x=8.4
x=120 lbs soybeans
280-120=160 lbs corn
Stay in school, have went is very incorrect, it should be have gone.
Have a good day.